Ana ebu...
|
Pịa ma Jide iji Dọkpụrụ Ya |
|||
|
Pịa Ebe a ka Imechi |
|||
Ajụjụ 2 Ripọtì
Ajụjụ 3 Ripọtì
Ajụjụ 4 Ripọtì
Ajụjụ 6 Ripọtì
Ajụjụ 7 Ripọtì
In this figure, \(PQ = PR = PS\) and \(\angle SRT = 68^\circ\). Find \(\angle QPS\).
i.e. 2(y + x) + QPS = 360∘
QPS = 360∘ - 2(y + x)
But x + y + 68∘ = 180∘
x + y = 180∘ - 68∘ = 180∘
x + y = 180∘ - 68∘
= 112∘
QPS = 360∘ - 2(112∘ )
360∘ - 224∘ = 136∘
Ajụjụ 8 Ripọtì
= 1.358
x 1001
= 33.752
= 16.875
= 1678
Ajụjụ 9 Ripọtì
Ajụjụ 10 Ripọtì
If \( \cos x = \sqrt{\frac{a}{b}} \), find coses x
Ajụjụ 11 Ripọtì
Rationalize \( \frac{2\sqrt{3}+3\sqrt{2}}{3\sqrt{2}-2\sqrt{3}} \)
Ajụjụ 12 Ripọtì
A car travels from calabar to Enugu, a distance of p km with an average speed of u km per hour and continues to benin, a distance of q km, with an average speed of w km per hour. Find its average speed from Calabar to Benin
Average speed = total DistanceTotal Time
from Calabar to Enugu in time t1, hence
t1 = pu
also from Enugu to Benin
t2 = qw
Average speed = p+qt1+t2
= p+qpu+qw
= p + q × uwpw+qu
= uw(p+q)pw+qu
Ajụjụ 13 Ripọtì
Simplify cos2 x (sec2x + sec2 x tan2x)
Ajụjụ 14 Ripọtì
If w varies inversely as \( \frac{ur}{u+r} \) and is equal to 8 when u = 2 and r = 6, find a relationship between u, v, w.
Ajụjụ 15 Ripọtì
PMN and PQR are two secants of the circle MQTRN and PT is a tangent. If PM = 5cm, PN = 12cm and PQ = 4.8cm, calculate the respective lengths of PR and PT in centimeters
Ajụjụ 16 Ripọtì
Ajụjụ 17 Ripọtì
Musa borrows N10.00 at 2% per month simple interest and repays N8.00 after 4 months. How much does he still owe?
I = PRT100
= 10×2×4100
= 45
= 0.8
Total amount = N10.80
He pays N8.00
Remainder = 10.80 - 8.00
= N2.80
Ajụjụ 18 Ripọtì
Ajụjụ 19 Ripọtì
Ajụjụ 20 Ripọtì
Ajụjụ 21 Ripọtì
Ajụjụ 22 Ripọtì
Ajụjụ 24 Ripọtì
Simplify \(2\log \frac{2}{5} - \log \frac{72}{125} + \log 9\)
Ajụjụ 25 Ripọtì
Find correct to 3 decimal places \(\left(\frac{1}{0.05} + \frac{1}{5.005}\right) - (0.05 x 2.05)\)
500.50.05
- 1025
100.1 - 0.1025
= 99.9975
99.998 to 3 sig. fiq.
Ajụjụ 26 Ripọtì
Simplify \(3\frac{1}{3} - 1\frac{1}{4} x \frac{2}{3} + 1\frac{2}{5}\)
3 - 13
- (54
x 23
) + 125
= 103
- 56
+ 75
= 100−25+4230
= 11730
= 3.9
≈
4
Ajụjụ 27 Ripọtì
If the exterior angles of a pentagon are xo, (x + 5)o, (x + 10), (x + 15)o and (x + 20)o, find x
Ajụjụ 28 Ripọtì
Ajụjụ 29 Ripọtì
Fifty boxes each of 50 bolts were inspected for the number which were defective. The following was the result:
| No. defective per box | 4 | 5 | 6 | 7 | 8 | 9 |
| No. of boxes | 2 | 7 | 17 | 10 | 8 | 6 |
The mean and the median of the distribution are respectively
Ajụjụ 30 Ripọtì
Fifty boxes each of 50 bolts were inspected for the number which were defective. The following was the result:
| No. defective per box | 4 | 5 | 6 | 7 | 8 | 9 |
| No. of boxes | 2 | 7 | 17 | 10 | 8 | 6 |
Find the percentage of boxes containing at least 5 defective bolts each.
Ajụjụ 32 Ripọtì
In the diagram above, |PQ| = |QR|, |PS| = |RS|, ∠PSR = 30o and ∠PQR = 80o. Find ∠SPQ.
Ajụjụ 33 Ripọtì
Simplify \( \frac{1}{1+\sqrt{5}} - \frac{1}{1-\sqrt{5}} \)
Ajụjụ 35 Ripọtì
Ajụjụ 36 Ripọtì
What is the product of \( \frac{27}{5^1}(3)^{-3} \) and \( \frac{(1)^{-1}}{5} \)?
Ajụjụ 38 Ripọtì
Ajụjụ 39 Ripọtì
A crate of soft drinks contains 10 bottles of Coca-cola, 8 of Fanta and 6 of sprite. If one bottle is selected at random, what is the probability that it is NOT a Cocacola bottle?
Coca-cola = 10 bottles, Fanta = 8 bottles
Sprite = 6 bottles, Total = 24
P(cola-cola) = 1024
P(not coca-cola) = 1 - 1024
24?1024
= 1424
= 724
Ajụjụ 40 Ripọtì
Ajụjụ 41 Ripọtì
Ajụjụ 43 Ripọtì
Ajụjụ 44 Ripọtì
Ajụjụ 46 Ripọtì
Evaluate \( \frac{xy^2-x^2y}{x^2-xy^1} \), when x = -2 and y = 2
xy2−x2yx2−xy1
= (−2)(3)2−(−2)2(3)(−2)2−(−2)(3)
= -3
Ajụjụ 47 Ripọtì
Ajụjụ 48 Ripọtì
Ị ga-achọ ịga n'ihu na omume a?